Ram tells shyam when I was half of your age you were one third of my present age. If ram's present age is 36. How old is shyam now?
Answer by lwsshak3(11628)

Ram tells shyam when I was half of your age you were one third of my present age. If ram's present age is 36. How old is shyam now?
***
let x=shyam's age now
12+(36-x/2)=x
48=1.5x
x=48/1.5=32
How old is shyam now? 32